OpenStudy (anonymous):

Find the roots : 1.) 5 sqrt x -x=6 2.) x^2 -8=17 please help !!

OpenStudy (raden):

1) 5sqrt(x) - x = 6 we knowed that x > 0 first, add by x to both sides, giving us 5sqrt(x) - x + x = 6 + x 5 sqrt(x) = 6 + x then square both sides, to get (5 sqrt(x))^2 = (x+6)^2 25x = x^2 + 12x + 36 simplify, now we have : x^2 - 13x + 36 = 0 solve for x

OpenStudy (raden):

this, can be factored : (x-4)(x-9) = 0 now, solve for x

OpenStudy (raden):

got it ? @Rea731

OpenStudy (anonymous):

Yes thank you I am just trying to transfer everything

OpenStudy (raden):

did u get number 2 ?